two planes leave an airport at the same time. one flies west at 600 mph and the other flies south at 800 mph. how far apart will they be in one hour?
Answer by checkley77(12569)

600^2+800^2=X^2
360,000+640,000=X^2
1,000,000=X^2
X=SQRT1,000,000
X=1,000 MILES IS THE DISTANCE BETWEEN THE 2 PLANES AFTER I HOUR.